TETSXPREME:

Learning with TET pt.3 Kindergarten Math ~

TETSXPREME:

1 attachment
jacksonsl:

LOL ok need help or some

MxxnLight:

You better pay attention. cause we aint doin tis for free.

jacksonsl:

exactly

MxxnLight:

B)

PapiTara:

1+1=? @TETSXPREME

jacksonsl:

do not give her the answer

nevaehhh:

93 i think 🌝

linahere:

tbh idk this either

MxxnLight:

94 is what I got-

TETSXPREME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @PapiTara 1+1=? @TETSXPREME \(\color{#0cbb34}{\text{End of Quote}}\) 2?

jacksonsl:

do not give her the Exact answer

PapiTara:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME \(\color{#0cbb34}{\text{Originally Posted by}}\) @PapiTara 1+1=? @TETSXPREME \(\color{#0cbb34}{\text{End of Quote}}\) 2? \(\color{#0cbb34}{\text{End of Quote}}\) 5?

MxxnLight:

💀✋🏻

linahere:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jacksonsl do not give her the Exact answer \(\color{#0cbb34}{\text{End of Quote}}\) him*

MxxnLight:

Dude we're not sure of the answer. ._.

sophia333:

192?

jacksonsl:

mk just sayin

nevaehhh:

48+24+12+6+3 72+18+3 90+3 93

sophia333:

:/

jacksonsl:

😀😀😀

MxxnLight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nevaehhh 48+24+12+6+3 72+18+3 90+3 93 \(\color{#0cbb34}{\text{End of Quote}}\) OHHH

nevaehhh:

wait im wrong-

MxxnLight:

You telle tubby. ._.

TETSXPREME:

so what is it? o;

nevaehhh:

48+24+24+12+12+6+6 48+48+24+12 96+36 132

jhonyy9:

- hits the floor 4 time first 48 to 24 2. 24 to 12 3 12 to 6 4 6 to 3 i think in this way is right @supie any idea pls here ?

jhonyy9:

ty .